# Denoising diffusion probabilistic models are optimally adaptive to unknown low dimensionality

### v2

@@ Abstract
The denoising diffusion probabilistic model (DDPM) has emerged as a mainstream generative model in generative AI. While sharp convergence guarantees have been established for the DDPM, the iteration complexity is, in general, proportional to the ambient data dimension, resulting in overly conservative theory that fails to explain its practical efficiency. This has motivated the recent work Li and Yan, 2024a to investigate how the DDPM can achieve sampling speed-ups through automatic exploitation of intrinsic low dimensionality of data. We strengthen this line of work by demonstrating, in some sense, optimal adaptivity to unknown low dimensionality. For a broad class of data distributions with intrinsic dimension $k$, we prove that the iteration complexity of the DDPM scales nearly linearly with $k$, which is optimal when using KL divergence to measure distributional discrepancy. Notably, our work is closely aligned with the independent concurrent work Potaptchik et al., (2024) — posted two weeks prior to ours — in establishing nearly linear-$k$ convergence guarantees for the DDPM.
@@ 1 Introduction
In generative modeling, one is asked to produce novel samples that resemble, in distribution, the data on which the model has been trained. In order to fully unleash its power, a generative modeling algorithm must exploit, either implicitly or explicitly, the distinctive features of the data distribution of interest, rather than merely optimizing for worst-case scenarios. A prominent example of such distinctive features is the low-dimensional structure underlying the data of interest, which permeates modern statistical modeling. For instance, image data often resides on some low-dimensional manifold (Simoncelli and Olshausen,, 2001; Pope et al.,, 2021), and genomics data might also be effectively modeled through manifolds (Moon et al.,, 2018; Zhu et al.,, 2018). While harnessing these low-dimensional data structures is generally believed to enable more efficient data generation, the theoretical underpinnings for this capability in contemporary generative modeling algorithms remain vastly under-explored.
@@
In this paper, we study how low-dimensional data structure influences the performance of diffusion models, an emerging paradigm that facilitates remarkable progress across various subfields of modern generative modeling (Chen et al., 2024a, ; Sohl-Dickstein et al.,, 2015; Dhariwal and Nichol,, 2021; Ho et al.,, 2022; Ramesh et al.,, 2022; Watson et al.,, 2023). While empirical evidence suggests that diffusion models might not suffer from the curse of dimensionality, a rigorious understanding of their efficacy in the presence of low-dimensional structure remains elusive. Consider, for concreteness, a mainstream diffusion model known as the denoising diffusion probabilistic model (DDPM) (Ho et al.,, 2020), which has received widespread adoption in practice and sparked considerable interest from the theoretical community to demystify its efficacy (e.g., Chen et al., 2022b ; Chen et al., 2022a ; Lee et al., (2022, 2023); Li et al., 2024b ; Tang, (2023); Mbacke and Rivasplata, (2023); Tang and Zhao, (2024); Liang et al., (2024); Li and Yan, 2024b ). In order for the DDPM sampler to generate a $d$-dimensional new sample that approximates well the target distribution in terms of KL divergence, the number of steps [^1] Given that each step of the DDPM requires evaluating a score function (typically through computing the output of a large neural network or transformer), the sampling speed of the DDPM is largely dictated by the number of steps. needs to scale linearly with $d$, a scaling that has been shown to be un-improvable in general (Benton et al.,, 2024). However, in application like image generation, the number of pixels in each image might oftentimes be exceedingly large, far exceeding the number of steps typically needed to run the DDPM in practice. This indicates that even linear scaling with $d$ does not adequately explain the observed efficiency of the DDPM in practice.
@@
To bridge the theory-practice gap for the data generation stage, the seminal work Li and Yan, 2024a demonstrated that DDPM can automatically adapt to the intrinsic dimension of the target distribution $p_{\mathsf{data}}$, which rigorously justifies sampling speed-ups if the support of $p_{\mathsf{data}}$ is intrinsically low-dimensional. As a particularly intriguing message of this prior result, the DDPM sampler, in its original form, achieves favorable adaptivity without explicitly modeling the low-dimensional structure.
@@
Despite this remarkable theoretical progress, however, it was unclear to what extent the low-dimensional structure could enhance the sampling efficiency of the DDPM. More precisely, Li and Yan, 2024a proved that when the support of $p_{\mathsf{data}}$ has intrinsic dimension $k$ (to be made precise in Section 4.1.1), the iteration complexity of the DDPM scales at most with $k^{4}$ (up to some logarithmic factor and other dependency on the target accuracy level), provided that KL divergence is used to measure the distributional discrepancy. This result was recently improved by Azangulov et al., (2024) to the order of $k^{3}$, assuming that data lies within some low-dimensional smooth manifolds. This raises a natural question: is the $k^{3}$ scaling essential, and if not, how to sharpen it? Addressing this question would provide a quantitative understanding about DDPM’s degree of adaptivity when tackling low-dimensional structure.
@@
In this work, we develop a sharp convergence theory for the DDPM in the face of unknown low dimensionality. For a broad family of target data distributions $p_{\mathsf{data}}$ with intrinsic dimension $k$, we show that it takes DDPM at most $O(k/\varepsilon^{2})$ (up to some logarithmic factor) steps to yield an $\varepsilon^{2}$-accurate distribution — measured by the KL divergence — assuming access to perfect score functions. This result, which shares similarity with a concurrent work Potaptchik et al., (2024) to be discussed momentarily, improves upon the iteration complexity derived by Azangulov et al., (2024) by a factor of $k^{2}$, and achieves nearly optimal scaling in $k$ without any burn-in requirement. Further, our convergence theory can be easily extended to accommodate imperfect score estimation.
To provide some intuition about its adaptivity, it is helpful to recognize that the original DDPM update rule by Ho et al., (2020) is precisely equivalent to running a carefully parameterized stochastic differential equation (SDE) upon discretization using the exponential integrator, a key observation that appeared in the recent work Azangulov et al., (2024). Crucially, the drift term of this discretized SDE is semi-linear, with the non-linear component proportional to the posterior mean of the data given its Gaussian corruption. Intuitively, when the support of $p_{\mathsf{data}}$ is low-dimensional, this nonlinear drift term acts as a “projection” onto the low-dimensional manifold of interest, harnessing this intrinsic structure to enhance the smoothness of the solution path and thereby speed up the sampling process. Somewhat surprisingly, while DDPM (Ho et al.,, 2020) was initially proposed to maximize a variational lower bound on the log-likelihood, the resulting update rule inherently adopts the desirable time discretization schedule, in a way that is fully adaptive to unknown low-dimensional distributions. A key ingredient underlying our analysis of the discretized SDE lies in carefully calibrating the evolution of the posterior mean function along the forward process.
The concurrent work Potaptchik et al., (2024).
While finalizing this work, we became aware of an interesting concurrent work Potaptchik et al., (2024), posted two weeks before our manuscript, that delivered the first convergence guarantees exhibiting linear dependency on the intrinsic manifold dimension of the data. Our paper, which stands as an independent and concurrent contribution, has similar aim and establishes similar results as Potaptchik et al., (2024), albeit under different assumptions on the low-dimensional structure. We shall provide more detailed comparisons between our work and Potaptchik et al., (2024) in Section 4.2.
@@ 2 Preliminaries
  • Data generation process. In order to operate in discrete time, let us choose a set of $N$ time discretization points $0=t_{0}<t_{1}<\dots<t_{N}<t_{N+1}=T$ for some large enough $T$, and adopt the update rule below: [^2] While the DDPM update rule was introduced in Ho et al., (2020) through the choices of $\{\alpha_{n}\}$ instead of the time discretization points $\{t_{n}\}$, it is equivalent to the version adopted herein.
@@
    which are fully determined by the locations of the time discretization points. In a nutshell, each iteration in (2) computes the weighted superposition of the current iterate, its corresponding score estimate, and an independent Gaussian vector, where the weights were originally proposed by Ho et al., (2020) to optimize a variational lower bound on the log-likelihood.
@@
The feasibility to reverse the forward process with the aid of score functions can be elucidated in the continuous-time limits, an approach popularized by Song et al., (2021). To be more specific, construct the forward process as the standard Ornstein-Uhlenbeck (OU) process (Oksendal and Oksendal,, 2003)
@@
where $(B_{t})$ represents another standard Brownian motion in $\mathbb{R}^{d}$ independent from $(X_{t})$ and $(W_{t})$. If $Y_{0}\sim X_{T}$, then classical results in the SDE literature (Anderson,, 1982; Haussmann and Pardoux,, 1986) tell us that
@@
holds under fairly mild regularity conditions on $p_{\mathsf{data}}$, thereby confirming the plausibility of score-based generative models. Note, however, that practical implementation of (5) has to operate in discrete time, and the sampling fidelity relies heavily upon the time discretization error. In fact, a large body of prior convergence theory for DDPM-type stochastic samplers (e.g., Chen et al., 2022b ; Chen et al., 2023a ; Benton et al., (2024)) was established by studying a certain time-discretized version of (5), with $s_{T-t}(\cdot)$ replaced by the score estimate $\widehat{s}_{T-t}(\cdot)$ to account for the effect of score estimation errors.
@@
one can readily derive from Tweedie’s formula (Efron,, 2011) that
@@ 3 DDPM as an adaptively discretized SDE
While previous studies (e.g., Chen et al., 2022b ; Chen et al., 2023a ; Benton et al., (2024)) have explored the convergence properties of time discretization of the ideal reverse process (5) through the SDE framework, a dominant fraction of these results fell short of directly addressing the precise update rule (2) of the DDPM sampler. As pointed out by Li and Yan, 2024a , the coefficients (or weights) adopted in the DDPM-type sampler are critical for effective adaptation to unknown low dimensionality; in fact, even mild perturbation of the coefficients can lead to considerable performance degradation, a curious phenomenon that does not occur in the general full-dimensional setting.
@@
After the initial posting of this paper, the authors of the concurrent work Potaptchik et al., (2024) informed us that a similar SDE already appeared in the recent work Azangulov et al., (2024, Section 7.2) as a type of backward SDE discretization with first-order score correction, where they discussed the motivation of score correction and its effect upon the discretization error. See also Azangulov et al., (2024, Page 90) for their justification of the SDE equivalence of the DDPM.
Importantly, solving the above SDE (11) is equivalent to running the original DDPM update rule (2), as asserted by the following theorem.
@@
$Y_{t_{n+1}}$ $=\frac{1}{\sqrt{\alpha_{n}}}\big{(}Y_{t_{n}}+(1-\alpha_{n})\widehat{s}_{T-t_{n}}(Y_{t_{n}})\big{)}+\sqrt{\frac{(1-\alpha_{n})(1-\overline{\alpha}_{n+1})}{1-\overline{\alpha}_{n}}}\widetilde{Z}_{n}$ (12)
for some random vectors $\{\widetilde{Z}_{n}\}$ in $\mathbb{R}^{d}$, where the $\widetilde{Z}_{n}$’s are statistically independent obeying $\widetilde{Z}_{n}\sim\mathcal{N}(0,I_{d})$.
With this result in place, the equivalence of the solution (12) and the original DDPM update rule (2) is self-evident. The proof of Theorem 1 is postponed to Appendix A.
Recall that in the general full-dimensional settings, previous work (e.g., Chen et al., 2023a ) suggested that discretizing different parametrization of (5) typically results in the same order of discretization errors, indicating that distinguishing between them might be unnecessary. This message, however, does not hold in the presence of low-dimensional structure, which merits further discussions as follows.
  • A reparameterized semi-linear SDE. As can be readily seen, if one has access to the true score function (i.e., $\widehat{s}_{t}(\cdot)=s_{t}(\cdot)$), then SDE (11b) can be viewed as the discretization of the true reverse SDE under the exponential integrator scheme:
@@
    where the last line invokes the Tweedie formula (8). In other words, the key is to rearrange the drift term of the reverse process (5) into a different semi-linear form, with the nonlinear component reflecting the posterior mean rather than the score.
@@
  • Parameterization matters. It is worth noting that how one re-parameterizes the drift term of (5) — namely, how one rearranges the semi-linear drift term therein — can exert significant influences upon the sampling efficiency. This can be elucidated via the inspiring example singled out in Li and Yan, 2024a (, Theorem 2). More concretely, when the target data distribution is isotropic Gaussian on a $k$-dimension linear subspace ($k\ll d$), the iteration complexity of the DDPM sampler is proportional to $O(\mathrm{poly}(k))$ (Li and Yan, 2024a, ), but most of other DDPM-type update rules (i.e., the ones with different coefficients than the original DDPM sampler) can take $\Omega(\mathrm{poly}(d))$ steps to yield the target sampling quality. Informally speaking, parameterizing the semi-linear drift term differently might result in a solution within $O(d/\mathrm{poly}(N))$ KL divergence from the DDPM solution [^3] Roughly speaking, in view of the Taylor expansion, we know that the discrepancy between different solutions is of order $O(1/\mathrm{poly}(N))$; but since $d$-dimensional Gaussian noise is added in each step of the sampling procedure, this discrepancy can result in $O(d/\mathrm{poly}(N))$ difference in KL divergence per step.; this discrepancy is negligible in the full-dimensional setting, but could blow up when $k$ is exceedingly small compared to $d$ and when $N$ is proportional to $k$. All this underscores the critical importance of identifying a parameterization that adapts naturally to the intrinsic data geometry.

### v3

@@ Abstract
The denoising diffusion probabilistic model (DDPM) has emerged as a mainstream generative model in generative AI. While sharp convergence guarantees have been established for the DDPM, the iteration complexity is, in general, proportional to the ambient data dimension, resulting in overly conservative theory that fails to explain its practical efficiency. This has motivated the recent work Li and Yan (2024a) to investigate how the DDPM can achieve sampling speed-ups through automatic exploitation of intrinsic low dimensionality of data. We strengthen this line of work by demonstrating, in some sense, optimal adaptivity to unknown low dimensionality. For a broad class of data distributions with intrinsic dimension $k$, we prove that the iteration complexity of the DDPM scales nearly linearly with $k$, which is optimal when using KL divergence to measure distributional discrepancy.
@@ 1 Introduction
In generative modeling, one is asked to produce novel samples that resemble, in distribution, the data on which the model has been trained. In order to fully unleash its power, a generative modeling algorithm must exploit, either implicitly or explicitly, the distinctive features of the data distribution of interest, rather than merely optimizing for worst-case scenarios. A prominent example of such distinctive features is the low-dimensional structure underlying the data of interest, which permeates modern statistical modeling. For instance, image data often resides on some low-dimensional manifold (Simoncelli and Olshausen, 2001; Pope et al., 2021), and genomics data might also be effectively modeled through manifolds (Moon et al., 2018; Zhu et al., 2018). In fact, the ImageNet dataset with ambient dimension $d\approx 10^{4}$ is estimated to have an intrinsic dimension of $k\approx 43$ (Pope et al., 2021), which is significantly smaller than the ambient dimension. While harnessing these low-dimensional data structures is generally believed to enable more efficient data generation, the theoretical underpinnings for this capability in contemporary generative modeling algorithms remain vastly under-explored.
@@
In this paper, we study how low-dimensional data structure influences the performance of diffusion models, an emerging paradigm that facilitates remarkable progress across various subfields of modern generative modeling (Chen et al., 2024a; Sohl-Dickstein et al., 2015; Dhariwal and Nichol, 2021; Ho et al., 2022; Ramesh et al., 2022; Watson et al., 2023). Empirical evidence suggests that diffusion models might not suffer from the curse of dimensionality. Some of the empirical methods explicitly exploit dimension reduction designs like latent diffusion to leverage low dimensionality (Rombach et al., 2022), while in the meantime vanilla diffusion models also seem to exhibit adaptivity to unknown low-dimensional structure in practice. However, a rigorous understanding of the efficacy of standard diffusion models in the presence of low-dimensional structure remains elusive. Consider, for concreteness, a mainstream diffusion model known as the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020), which has received widespread adoption in practice and sparked considerable interest from the theoretical community to demystify its efficacy (e.g., Chen et al. (2022b, a); Lee et al. (2022, 2023); Li et al. (2025b); Tang (2023); Mbacke and Rivasplata (2023); Tang and Zhao (2024); Li et al. (2025a); Liang et al. (2024); Li and Yan (2024b)). In order for the DDPM sampler to generate a $d$-dimensional new sample that approximates well the target distribution in terms of KL divergence, the number of steps [^1] Given that each step of the DDPM requires evaluating a score function (typically through computing the output of a large neural network or transformer), the sampling speed of the DDPM is largely dictated by the number of steps. needs to scale linearly with $d$, a scaling that has been shown to be un-improvable in general (Benton et al., 2024). However, in application like image generation, the number of pixels in each image might oftentimes be exceedingly large, far exceeding the number of steps typically needed to run the DDPM in practice. This indicates that even linear scaling with $d$ does not adequately explain the observed efficiency of the DDPM in practice.
@@
To bridge the theory-practice gap for the data generation stage, the seminal work by Li and Yan (2024a) demonstrated that DDPM can automatically adapt to the intrinsic dimension of the target distribution $p_{\mathsf{data}}$, which rigorously justifies sampling speed-ups if the support of $p_{\mathsf{data}}$ is intrinsically low-dimensional. As a particularly intriguing message of this prior result, the DDPM sampler, in its original form, achieves favorable adaptivity without explicitly modeling the low-dimensional structure.
@@
Despite this remarkable theoretical progress, however, it was unclear to what extent the low-dimensional structure could enhance the sampling efficiency of the DDPM. More precisely, Li and Yan (2024a) proved that when the support of $p_{\mathsf{data}}$ has intrinsic dimension $k$ (to be made precise in Section 4.1.1), the iteration complexity of the DDPM scales at most with $k^{4}$ (up to some logarithmic factor and other dependency on the target accuracy level), provided that KL divergence is used to measure the distributional discrepancy. This result was recently improved by Azangulov et al. (2024) to the order of $k^{3}$, assuming that data lies within some low-dimensional smooth manifolds. This raises a natural question: is the $k^{3}$ scaling essential, and if not, how to sharpen it? Addressing this question would provide a quantitative understanding about DDPM’s degree of adaptivity when tackling low-dimensional structure.
@@
In this work, we develop a sharp convergence theory for the DDPM in the face of unknown low dimensionality. For a broad family of target data distributions $p_{\mathsf{data}}$ with intrinsic dimension $k$, we show that it takes DDPM at most $O(k/\varepsilon^{2})$ (up to some logarithmic factor) steps to yield an $\varepsilon^{2}$-accurate distribution — measured by the KL divergence — assuming access to perfect score functions. This result, which shares similarity with a concurrent work by Potaptchik et al. (2024) to be discussed momentarily, improves upon the iteration complexity derived by Azangulov et al. (2024) by a factor of $k^{2}$, and achieves nearly optimal scaling in $k$ without any burn-in requirement. Further, our convergence theory can be easily extended to accommodate imperfect score estimation.
To provide some intuition about its adaptivity, it is helpful to recognize that the original DDPM update rule by Ho et al. (2020) is precisely equivalent to running a carefully parameterized stochastic differential equation (SDE) upon discretization using the exponential integrator, a key observation that appeared in the recent work by Azangulov et al. (2024). Crucially, the drift term of this discretized SDE is semi-linear, with the non-linear component proportional to the posterior mean of the data given its Gaussian corruption. Intuitively, when the support of $p_{\mathsf{data}}$ is low-dimensional, this nonlinear drift term acts as a “projection” onto the low-dimensional manifold of interest, harnessing this intrinsic structure to enhance the smoothness of the solution path and thereby speed up the sampling process. Somewhat surprisingly, while DDPM was initially proposed to maximize a variational lower bound on the log-likelihood (Ho et al., 2020), the resulting update rule inherently adopts the desirable time discretization schedule, in a way that is fully adaptive to unknown low-dimensional distributions. A key ingredient underlying our analysis of the discretized SDE lies in carefully calibrating the evolution of the posterior mean function along the forward process.
The concurrent work Potaptchik et al. (2024).
While finalizing this work, we became aware of an interesting concurrent work Potaptchik et al. (2024), posted two weeks before our manuscript, that delivered the first convergence guarantees exhibiting linear dependency on the intrinsic manifold dimension of the data. Our paper, which stands as an independent and concurrent contribution, has similar aim and establishes similar results as Potaptchik et al. (2024), albeit under different assumptions on the low-dimensional structure. We shall provide more detailed comparisons between our work and Potaptchik et al. (2024) in Section 4.2.
@@ 2 Preliminaries
  • Data generation process. In order to operate in discrete time, let us choose a set of $N$ time discretization points $0=t_{0}<t_{1}<\dots<t_{N}<t_{N+1}=T$ for some large enough $T$, and adopt the update rule below: [^2] While the DDPM update rule was introduced in Ho et al. (2020) through the choices of $\{\alpha_{n}\}$ instead of the time discretization points $\{t_{n}\}$, it is equivalent to the version adopted herein.
@@
    which are fully determined by the locations of the time discretization points. In a nutshell, each iteration in (2) computes the weighted superposition of the current iterate, its corresponding score estimate, and an independent Gaussian vector, where the weights were originally proposed by Ho et al. (2020) to optimize a variational lower bound on the log-likelihood.
@@
The feasibility to reverse the forward process with the aid of score functions can be elucidated in the continuous-time limits, an approach popularized by Song et al. (2021). To be more specific, construct the forward process as the standard Ornstein-Uhlenbeck (OU) process (Oksendal and Oksendal, 2003)
@@
where $(B_{t})$ represents another standard Brownian motion in $\mathbb{R}^{d}$ independent from $(X_{t})$ and $(W_{t})$. If $Y_{0}\sim X_{T}$, then classical results in the SDE literature (Anderson, 1982; Haussmann and Pardoux, 1986) tell us that
@@
holds under fairly mild regularity conditions on $p_{\mathsf{data}}$, thereby confirming the plausibility of score-based generative models. Note, however, that practical implementation of (5) has to operate in discrete time, and the sampling fidelity relies heavily upon the time discretization error. In fact, a large body of prior convergence theory for DDPM-type stochastic samplers (e.g., Chen et al. (2022b, 2023a); Benton et al. (2024)) was established by studying a certain time-discretized version of (5), with $s_{T-t}(\cdot)$ replaced by the score estimate $\widehat{s}_{T-t}(\cdot)$ to account for the effect of score estimation errors.
@@
one can readily derive from Tweedie’s formula (Efron, 2011) that
@@ 3 DDPM as an adaptively discretized SDE
While previous studies (e.g., Chen et al. (2022b, 2023a); Benton et al. (2024)) have explored the convergence properties of time discretization of the ideal reverse process (5) through the SDE framework, a dominant fraction of these results fell short of directly addressing the precise update rule (2) of the DDPM sampler. As pointed out by Li and Yan (2024a), the coefficients (or weights) adopted in the DDPM-type sampler are critical for effective adaptation to unknown low dimensionality; in fact, even mild perturbation of the coefficients can lead to considerable performance degradation, a curious phenomenon that does not occur in the general full-dimensional setting.
@@
After the initial posting of this paper, the authors of the concurrent work Potaptchik et al. (2024) informed us that a similar SDE already appeared in the recent work (Azangulov et al., 2024, Section 7.2) as a type of backward SDE discretization with first-order score correction, where they discussed the motivation of score correction and its effect upon the discretization error. See also Azangulov et al. (2024, Page 90) for their justification of the SDE equivalence of the DDPM.
Importantly, SDE (11) is a continuous interpolation of the DDPM update rule (2), which means that solving the above SDE (11) leads to exactly the original DDPM sampler, as asserted by the following theorem.
@@
$Y_{t_{n+1}}$ $=\frac{1}{\sqrt{\alpha_{n}}}\big(Y_{t_{n}}+(1-\alpha_{n})\widehat{s}_{T-t_{n}}(Y_{t_{n}})\big)+\sqrt{\frac{(1-\alpha_{n})(1-\overline{\alpha}_{n+1})}{1-\overline{\alpha}_{n}}}\widetilde{Z}_{n}$
for statistically independent random vectors $\{\widetilde{Z}_{n}\}$ in $\mathbb{R}^{d}$ obeying $\widetilde{Z}_{n}\sim\mathcal{N}(0,I_{d})$, which recovers the DDPM sampler (2).
The proof of Theorem 1 is postponed to Appendix A.
Recall that in the general full-dimensional settings, previous work (e.g., Chen et al. (2023a)) suggested that discretizing different parametrization of (5) typically results in the same order of discretization errors, indicating that distinguishing between them might be unnecessary. This message, however, does not hold in the presence of low-dimensional structure, which merits further discussions as follows.
  • A reparameterized semi-linear SDE. As can be readily seen, if one has access to the true score function (i.e., $\widehat{s}_{t}(\cdot)=s_{t}(\cdot)$), then for some $\delta>0$, SDE (11b) can be viewed as the discretization of the true reverse SDE in the time horizon $0,T-\delta]$ under the exponential integrator scheme:
@@
    where the last line invokes the Tweedie formula (8). Here, the stopping time $\delta$ is introduced to avoid the singularity of the drift term when $t$ approaches $T$. In other words, the key is to rearrange the drift term of the reverse process (5) into a different semi-linear form, with the nonlinear component reflecting the posterior mean rather than the score.
@@
  • Parameterization matters. It is worth noting that how one re-parameterizes the drift term of (5) — namely, how one rearranges the semi-linear drift term therein — can exert significant influences upon the sampling efficiency. This can be elucidated via the inspiring example singled out in Li and Yan (2024a, Theorem 2). More concretely, when the target data distribution is isotropic Gaussian on a $k$-dimension linear subspace ($k\ll d$), the iteration complexity of the DDPM sampler (2) is proportional to $O(\mathrm{poly}(k))$ (Li and Yan, 2024a), but most of other DDPM-type update rules (i.e., the ones with different coefficients than the original DDPM sampler) can take $\Omega(\mathrm{poly}(d))$ steps to yield the target sampling quality. Informally speaking, parameterizing the semi-linear drift term differently might result in a solution within $O(d/\mathrm{poly}(N))$ KL divergence from the DDPM solution.  [^3] Roughly speaking, in view of the Taylor expansion, we know that the discrepancy between different solutions is of order $O(1/\mathrm{poly}(N))$; but since $d$-dimensional Gaussian noise is added in each step of the sampling procedure, this discrepancy can result in $O(d/\mathrm{poly}(N))$ difference in KL divergence per step. Note that this discrepancy is negligible in the full-dimensional setting, but could dominate the sampling error when $k$ is exceedingly small compared to $d$ and when $N$ is approximately proportional to kk. [^4] This order of $N$ coincides with the number of iterations we suggest in Corollary 1 for the DDPM sampler (up to logarithmic factor). All this underscores the critical importance of identifying a parameterization that adapts naturally to the intrinsic data geometry.